# Beam tracing and profile evolution for localised beams in inhomogeneous plasmas

### v2

@@ II.2 The beam tracing equations
$D_{ij}\hat{e}_{j}=H\hat{e}_{i}=0,$ (15)
where we have reintroduced $\frac{c}{\Omega}$ to make $D_{ij}$ dimensionless:
$D_{ij}(\bm{K},\bm{r})=\frac{c^{2}}{\Omega^{2}}(\delta_{il}\delta_{jm}-\delta_{ij}\delta_{ml})K_{m}K_{l}+\epsilon_{ij}.$ (16)
@@
$\begin{split}\partial K_{i}H&=\frac{\mathrm{d}q_{i}}{\mathrm{d}\tau},\\ \partial_{i}H&=-\frac{\mathrm{d}K_{i}}{\mathrm{d}\tau}.\end{split}$ (17)
At second order, we obtain a second order PDE for the evolution of the beam, which is the eikonal equation. We firstly obtain the beam envelope evolution equation by simply defining it:
$\frac{\mathrm{d}\bm{\Psi}}{\mathrm{d}\tau}+\bm{\Psi}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}(H)\bm{\Psi}+\bm{\Psi}\bm{\nabla}_{\bm{K}}\bm{\nabla}(H)+\bm{\nabla}\bm{\nabla}_{\bm{K}}(H)\bm{\Psi}+\bm{\nabla}\bm{\nabla}(H)=0,$ (18)
where the second form of the equation involves the $\tau$ derivative on the mathematical object $\bm{\Psi}$, thus it does not use the covariant definition. Next are the amplitude evolution equations, where $\alpha^{(0)}=|\alpha^{(0)}|\exp(\mathrm{i}\phi)$. We also define them such that we have the following equation for the magnitude of our amplitude:
$|\alpha^{(0)}|=C\frac{\mathrm{Det(}\Im(\bm{\Psi}))^{\frac{1}{4}}}{g^{\frac{1}{2}}},$ (19)
and the phase of our amplitude:
$\frac{\mathrm{d}\phi}{\mathrm{d}\tau}=\mathrm{i}\hat{e}^{*}_{i}\frac{\mathrm{d}\hat{e}_{i}}{\mathrm{d}\tau}+\frac{\mathrm{i}}{2}(\partial_{\mu}\hat{e}^{*}_{i}D_{im}\partial K_{\mu}\hat{e}_{m}-\partial K_{\mu}\hat{e}^{*}_{i}D_{im}\partial_{\mu}\hat{e}_{m})-\frac{1}{2}\Im(\Psi_{\mu\nu})\partial K_{\mu}\partial K_{\nu}H.$ (20)
@@
$\frac{\mathrm{d}\phi_{P}}{\mathrm{d}\tau}=\mathrm{i}\hat{e}^{*}_{i}\frac{\mathrm{d}\hat{e}_{i}}{\mathrm{d}\tau}+\frac{\mathrm{i}}{2}(\partial_{\mu}\hat{e}^{*}_{i}D_{im}\partial K_{\mu}\hat{e}_{m}-\partial K_{\mu}\hat{e}^{*}_{i}D_{im}\partial_{\mu}\hat{e}_{m}),$ (21)
@@
$\frac{\mathrm{d}\phi_{G}}{\mathrm{d}\tau}=-\frac{1}{2}\Im(\Psi_{\mu\nu})\partial K_{\mu}\partial K_{\nu}H.$ (22)
All these definitions reduce our second-order PDE into this relatively much simpler and crucial equation defining the evolution of our beam profile:
$\frac{\partial\mathcal{P}}{\partial\tau}-\mathrm{i}\left(\frac{1}{2}\partial K_{\mu}\partial K_{\nu}H\partial_{\mu}\partial_{\nu}(\mathcal{P})^{(0)}\right)+w_{\mu}T_{\mu\nu}\partial_{\nu}\mathcal{P}^{(0)}=0.$ (23)
where $T_{\mu\nu}$ is given by:
$\begin{split}T_{\mu\nu}(\tau)&=\left(\frac{\mathrm{d}\hat{\bm{y}}}{\mathrm{d}\tau}\cdot\hat{\bm{x}}\right)(\hat{x}_{\mu}\hat{y}_{\nu}-\hat{y}_{\mu}\hat{x}_{\nu})+\Big(\Psi_{\mu\rho}\partial K_{\rho}\partial K_{\nu}H+\partial_{\mu}\partial K_{\nu}H\Big)\end{split}$ (24)
This is a Fokker-Planck equationRisken1996, and is quite similar to the Ornstein-Uhlenbeck equationVatiwutipong2019. Whilst this is a great simplification from the Helmholtz equation(a 3D non-linear PDE), it still is not the most computationally efficient equation to solve.
All of these equations are ODEs which can be simply integrated over, with the exception of our profile evolution equation. While we have efficient methods for solving ODEs, solving a 2D linear PDE is still computationally more demanding. The primary purpose of our work is to be able to provide real-time beam analysis, thus possessing a faster approach would be desirable. In the subsequent section, we will do just that, by solving for generic families of solutions to eq. 23, by repeated application of independent ladder operators.
@@ III Ladder operator solutions of profile evolution equation
In this section, we present general operations that could yield the evolution of any initially confined beam profile, via ladder operators. When looking for solutions to eq. 23, we are finding eigenfunctions of the differential operator $\hat{\mathcal{D}}$ given by:
$\frac{\partial}{\partial\tau}-\mathrm{i}\frac{1}{2}\partial K_{\mu}\partial K_{\nu}H\partial_{\mu}\partial_{\nu}+w_{\mu}T_{\mu\nu}\partial_{\nu},$ (25)
where:
@@
The choice of 0 eigenvalue is always possible, since we are free to arbitrarily scale any eigenfunctions by a factor $\beta(\tau)$, which can be used to remove any non-zero eigenvalues. Although this approach gives us the approach to model the evolution of a great many different initial profiles, our end goal is to construct eigenfunctions that reduce to Gauss-Hermite beams at the plasma boundary. If we could do so, we can utilise the orthonormality of the Hermite polynomials to decompose any function into a sum of Gauss-Hermite modes, each of which we can evolve exactly. This avoids the need for cumbersome numerical simulation of eq. 23. Furthermore, the Gauss-Hermite solutions have been studied before, allowing us to discuss the validity of previous results in section V.2.
We will first construct solutions using generic ladder operators, and then see what specific choice of ladder operators will yield our chosen basis at the plasma boundary.
Before constructing said ladder operators, it is quite instructive to analyse what happens if one inputs a Gaussian profile as our beam profile. This is not only an important sanity check for eq. 23, but it also serves a crucial algebraic purpose later in our construction of the ladder operators.
@@ III.1 Gaussian profiles
As we pointed out, one is free to decompose the profile in any way you like. Thus, even if our beam was Gaussian, we could pull out a different Gaussian envelope, with the profile being the rest of our Gaussian beam(our profile is therefore Gaussian). We then expect our profile evolution equation to make our Gaussian profile evolve identically to the beam-tracing equations. It turns out that it does evolve exactly like the beam-tracing equations, so our model passes this basic sanity check. We outline the details of this derivation below.
@@
$\mathcal{P}=A(\tau)\exp\left(\frac{\mathrm{i}}{2}\bm{w}\cdot\bm{\Phi}(\tau)\cdot\bm{w}\right),$ (27)
and we have that:
@@
As pointed out during our derivation, we define $\bm{\Psi}$ to obey eq. 18. For consistency’s sake, we also expect $\bm{\Psi}_{\mathrm{tot}}$ to obey the same equation. Substituting the expression for our profile into eq. 23, we will have terms proportional to $\bm{ww}$, and terms that are independent. The terms independent of $\bm{w}$ are:
$\frac{\mathrm{d}\ln(A(\tau))}{\mathrm{d}\tau}+\frac{1}{2}\Phi_{\mu\nu}\partial K_{\mu}\partial K_{\nu}H=0.$ (29)
The expression for the terms proportional to $\bm{ww}$ is much longer, thus, we will utilise matrix convention for brevity. These terms are:
$\begin{split}\frac{\mathrm{i}}{2}\bm{ww}:\bigg(&\frac{\mathrm{d}\bm{\Phi}}{\mathrm{d}\tau}+\bm{\Phi}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\bm{\Phi}+(\bm{\Phi}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\bm{\Psi}+\bm{\Psi}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\bm{\Phi})\\ &+(\bm{\Phi}\bm{\nabla}_{\bm{K}}\bm{\nabla}H+\bm{\nabla}\bm{\nabla}_{\bm{K}}H\bm{\Phi})\bigg)=0.\end{split}$ (30)
Our end goal is to show that this obeys the beam-tracing equations. It is easier to get to here from the beam-tracing equations than it is to go the other way round, so we shall begin from there. Note that $\bm{\Psi}$ obeys eq. 18, as would $\bm{\Psi}_{\mathrm{tot}}$, which is what we are setting out to prove. Assuming that it does, we will see if we arrive at a contradiction. Assuming that the following is true:
$\begin{split}&\frac{\mathrm{d}(\bm{\Psi}+\bm{\Phi})}{\mathrm{d}\tau}+(\bm{\Psi}+\bm{\Phi})\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H(\bm{\Psi}+\bm{\Phi})\\ &+\bm{\nabla}\bm{\nabla}_{\bm{K}}H(\bm{\Psi}+\bm{\Phi})+(\bm{\Psi}+\bm{\Phi})\bm{\nabla}_{\bm{K}}\bm{\nabla}H+\bm{\nabla}\bm{\nabla}H=0,\end{split}$ (31)
utilising eq. 18, reduces this to:
$\begin{split}&\frac{\mathrm{d}\bm{\Phi}}{\mathrm{d}\tau}+\bm{\Phi}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\bm{\Phi}\\ &+\bm{\Psi}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\bm{\Phi}+\bm{\Phi}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\bm{\Psi}\\ &+\bm{\nabla}\bm{\nabla}_{\bm{K}}H\bm{\Phi}+\bm{\Phi}\bm{\nabla}_{\bm{K}}\bm{\nabla}H=0.\end{split}$ (32)
@@
This is precisely what our terms dependent on $\bm{x}\bm{x}$ tell us, albeit projected onto the $\bm{ww}$-[plane. This therefore confirms that if we were to have a Gaussian profile, its envelope must obey eq. 139, which is the same as obeying eq. 18(we obtained this by adding terms that equal 0). We can now turn our attention to the terms independent of $\bm{w}$. Note that if we add eq. 143 to our terms, we end up with:
$\frac{\mathrm{dln}(A(\tau)\alpha^{(0)}(\tau))}{\mathrm{d}\tau}+\frac{\mathrm{d}K_{\mu}}{\mathrm{d}\tau}\hat{e}^{*}_{m}\partial K_{\mu}\hat{e}_{m}+\hat{e}^{*}_{i}\partial K_{\mu}D_{im}\partial_{\mu}\hat{e}_{m}+\frac{1}{2}(\Psi_{\mu\nu}+S_{\mu\rho}\Phi_{\rho\sigma}S_{\rho\nu})\partial K_{\mu}\partial K_{\nu}H=0.$ (33)

### v3

@@ II.2 The beam tracing equations
$\bm{D}\cdot\hat{\bm{e}}=H\hat{\bm{e}}=0,$ (16)
where:
$\bm{D}(\bm{K},\bm{r})=\frac{c^{2}}{\Omega^{2}}(\bm{KK}-K^{2}\bm{1})+\bm{\epsilon}.$ (17)
@@
$\begin{split}\bm{\nabla}_{\bm{K}}H&=\frac{\mathrm{d}\bm{q}}{\mathrm{d}\tau},\\ \bm{\nabla}H&=-\frac{\mathrm{d}\bm{K}}{\mathrm{d}\tau}.\end{split}$ (18)
At second order, we obtain a second order PDE for the evolution of the beam, which is the eikonal equation. We firstly obtain the beam envelope evolution equation by simply defining it to be:
$\frac{\mathrm{d}\bm{\Psi}}{\mathrm{d}\tau}+\bm{\Psi}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}(H)\cdot\bm{\Psi}+\bm{\Psi}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}(H)+\bm{\nabla}\bm{\nabla}_{\bm{K}}(H)\cdot\bm{\Psi}+\bm{\nabla}\bm{\nabla}(H)=0,$ (19)
Next are the amplitude evolution equations, where $A^{(0)}=|A^{(0)}|\exp(\mathrm{i}\phi)$. We also define them such that we have the following equation for the magnitude of the complex amplitude:
$|A^{(0)}|=C\frac{\mathrm{Det(}\Im(\bm{\Psi}))^{\frac{1}{4}}}{g^{\frac{1}{2}}},$ (20)
and the phase of the complex amplitude:
$\frac{\mathrm{d}\phi}{\mathrm{d}\tau}=\mathrm{i}\hat{\bm{e}}^{*}\cdot\frac{\mathrm{d}\hat{\bm{e}}}{\mathrm{d}\tau}+\frac{\mathrm{i}}{2}\left(\frac{\partial\hat{\bm{e}}^{*}}{\partial r_{\mu}}\cdot\bm{D}\cdot\frac{\partial\hat{\bm{e}}}{\partial K_{\mu}}-\frac{\partial\hat{\bm{e}}^{*}}{\partial K_{\mu}}\cdot\bm{D}\cdot\frac{\partial\hat{\bm{e}}}{\partial r_{\mu}}\right)-\frac{1}{2}\Im(\bm{\Psi}):\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H.$ (21)
@@
$\frac{\mathrm{d}\phi_{P}}{\mathrm{d}\tau}=\mathrm{i}\hat{\bm{e}}^{*}\cdot\frac{\mathrm{d}\hat{\bm{e}}}{\mathrm{d}\tau}+\frac{\mathrm{i}}{2}\left(\frac{\partial\hat{\bm{e}}^{*}}{\partial r_{\mu}}\cdot\bm{D}\cdot\frac{\partial\hat{\bm{e}}}{\partial K_{\mu}}-\frac{\partial\hat{\bm{e}}^{*}}{\partial K_{\mu}}\cdot\bm{D}\cdot\frac{\partial\hat{\bm{e}}}{\partial r_{\mu}}\right),$ (22)
@@
$\frac{\mathrm{d}\phi_{G}}{\mathrm{d}\tau}=-\frac{1}{2}\Im(\bm{\Psi}):\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H.$ (23)
All these definitions reduce our second-order PDE into this much simpler equation defining the evolution of our beam profile:
$\frac{\partial\mathcal{P}}{\partial\tau}-\mathrm{i}\left(\frac{1}{2}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H:\bm{\nabla}_{w}\bm{\nabla}_{w}\mathcal{P}\right)+\bm{w}\cdot\bm{T}_{w}\cdot\bm{\nabla}_{w}\mathcal{P}=0.$ (24)
where $\bm{T}$ is given by:
$\begin{split}\bm{T}(\tau)&=\left(\frac{\mathrm{d}\hat{\bm{y}}}{\mathrm{d}\tau}\cdot\hat{\bm{x}}\right)(\hat{\bm{x}}\hat{\bm{y}}-\hat{\bm{y}}\hat{\bm{x}})+\Big(\bm{\Psi}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H+\bm{\nabla}\bm{\nabla}_{\bm{K}}H\Big),\end{split}$ (25)
[…]
All of these equations are ODEs which can be simply integrated over, with the exception of our profile evolution equation. While we have efficient methods for solving ODEs, solving a 2D linear PDE is still computationally more demanding. The primary purpose of our work is to be able to provide real-time beam analysis, thus possessing a faster approach would be desirable. In the subsequent section, we will do just that, by solving for generic families of solutions to eq. 24 via repeated application of ladder operators.
@@ III Ladder operator solutions of profile evolution equation
In this section, we present general operations that could yield the evolution of any initially confined beam profile via ladder operators[Leen2016]. We are clearly interested solutions to eq. 24, so we are finding eigenfunctions of the differential operator $\hat{\mathcal{D}}$ given in index notation by:
$\hat{\mathcal{D}}=\frac{\partial}{\partial\tau}-\mathrm{i}\frac{1}{2}\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}H\frac{\partial}{\partial w_{\mu}}\frac{\partial}{\partial w_{\nu}}+w_{\mu}T_{\mu\nu}\frac{\partial}{\partial w_{\nu}},$ (26)
where $\frac{\partial}{\partial w_{\mu}}:=\nabla_{w,\mu}$ such that:
@@
Although this approach gives us the freedom to model the evolution of a multitude of initial profiles, our end goal will be to construct solutions that reduce to Gauss-Hermite beams at the plasma-vacuum boundary. If we could do so, we can utilise the orthonormality of the Hermite polynomials to decompose any function into a sum of Gauss-Hermite modes, each of which we can evolve exactly. Furthermore, the Gauss-Hermite solutions have been studied before, allowing us to discuss the consistency of previous results in section VI.
Since $\hat{\mathcal{D}}$ is a second order differential operator, the complete set of solutions to $\hat{\mathcal{D}}$ can be decomposed into two separate families of solutions. The basic idea is to construct two separate families of operators, that we call $\hat{\mathcal{L}}$ and $\hat{\mathcal{R}}$ [^2] This stands for left and right, for lack of a better description of the operators. This is due to the fact that they will not generally function as exact inverses or adjoints of each other, which is demonstrated in appendix B., that commute with $\hat{\mathcal{D}}$. That is, $\hat{\mathcal{D}},\hat{\mathcal{L}}]=\hat{\mathcal{D}}\hat{\mathcal{L}}-\hat{\mathcal{L}}\hat{\mathcal{D}}=0$, and similarly $\hat{\mathcal{D}},\hat{\mathcal{R}}]=0$. Within this paragraph, $\hat{\mathcal{L}}$ and $\hat{\mathcal{R}}$ each refer to any operator from each family of operators, the elements of which will be elucidated in eq. 46. These operators become useful if we know at least one particular solution $f$ to $\hat{\mathcal{D}}$. Using the commutativity with $\hat{\mathcal{L}}$ and $\hat{\mathcal{R}}$, this automatically implies that $\hat{\mathcal{L}}f$ and $\hat{\mathcal{R}}f$ are also solutions to $\hat{\mathcal{D}}$. We use this property to generate two families of solutions. which are given by $\hat{\mathcal{L}}^{n}f$ and $\hat{\mathcal{R}}^{n}f$. [^3] If $f$ is in the kernel of $\hat{\mathcal{L}}$ (or respectively $\hat{\mathcal{R}}$), we will only be able to generate the family corresponding to $\hat{\mathcal{R}}$ (respectively $\hat{\mathcal{L}}$). In that case, it will be necessary to find a second particular solution $f^{\prime}$ to $\hat{\mathcal{D}}$, which will yield the second family of solutions. For our current introductory purposes, we only present other important properties of our ladder operators in appendix B, where they are introduced in a more natural manner.
Before constructing said ladder operators, it will be useful to consider the following corollary. Suppose that we define $\bm{\Psi}^{\prime},\bm{\Psi},\bm{\Phi}$ to be complex-symmetric matrices obeying the following relation:
$\bm{\Psi}^{\prime}=\bm{\Psi}+\bm{\Phi}.$ (28)
If we constrain both $\bm{\Psi}^{\prime}$ and $\bm{\Psi}$ to obey eq. 19, then the following must be true:
$\begin{split}&\frac{\mathrm{d}(\bm{\Psi}+\bm{\Phi})}{\mathrm{d}\tau}+(\bm{\Psi}+\bm{\Phi})\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\cdot(\bm{\Psi}+\bm{\Phi})\\ &+\bm{\nabla}\bm{\nabla}_{\bm{K}}H\cdot(\bm{\Psi}+\bm{\Phi})+(\bm{\Psi}+\bm{\Phi})\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}H+\bm{\nabla}\bm{\nabla}H=0.\end{split}$ (29)
Utilising eq. 19, this time for $\bm{\Psi}$, reduces this to:
$\begin{split}&\frac{\mathrm{d}\bm{\Phi}}{\mathrm{d}\tau}+\bm{\Phi}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\cdot\bm{\Phi}\\ &+\bm{\Psi}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\cdot\bm{\Phi}+\bm{\Phi}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\cdot\bm{\Psi}\\ &+\bm{\nabla}\bm{\nabla}_{\bm{K}}H\cdot\bm{\Phi}+\bm{\Phi}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}H=0.\end{split}$ (30)
This is equivalent to saying that if we have symmetric matrices $\bm{\Psi},\bm{\Phi}$ obeying eq. 19 and eq. 30 respectively, we know that their sum $\bm{\Psi}+\bm{\Phi}$ must also obey eq. 19. This will turn out to be a very useful property in subsequent derivations.
It is first instructive to analyse what happens if one inputs a Gaussian profile as our beam profile. This is not only an important consistency check for eq. 24, but it also serves a crucial algebraic purpose later in our construction of the profiles generated by our ladder operators.
@@ III.1 Gaussian profiles
As we pointed out, one is free to decompose the profile in any way you like. Thus, even if our beam was exactly Gaussian, we could pull out a different Gaussian envelope, with the profile being the rest of our Gaussian beam(our profile is therefore Gaussian). We then expect our profile evolution equation to make our Gaussian profile evolve identically to the beam-tracing equations, which we subsequently prove.
@@
$\mathcal{P}=\alpha(\tau)\exp\left(\frac{\mathrm{i}}{2}\bm{w}\cdot\bm{\Phi}(\tau)\cdot\bm{w}\right),$ (31)
Substituting the expression for our profile into eq. 24, we will have terms proportional to $\bm{ww}$, and terms that are independent. The terms independent of $\bm{w}$ are:
$\frac{\mathrm{d}\ln(\alpha(\tau))}{\mathrm{d}\tau}+\frac{1}{2}\bm{\Phi}:\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H=0.$ (32)
The terms proportional to $\bm{ww}$ are given by:
$\begin{split}\frac{\mathrm{i}}{2}\bm{ww}:\bigg(&\frac{\mathrm{d}\bm{\Phi}}{\mathrm{d}\tau}+\bm{\Phi}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\cdot\bm{\Phi}+(\bm{\Phi}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\cdot\bm{\Psi}+\bm{\Psi}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\cdot\bm{\Phi})\\ &+(\bm{\Phi}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}H+\bm{\nabla}\bm{\nabla}_{\bm{K}}H\cdot\bm{\Phi})\bigg)=0.\end{split}$ (33)
With this choice for $\mathcal{P}$, we see from eq. 2 and eq. 3 that the total Gaussian envelope of the electric field, $\bm{\Psi}_{\mathrm{tot}}$, is given by the sum of the Gaussian envelope $\bm{\Psi}$, and the Gaussian envelope $\bm{\Phi}$ from $\mathcal{P}$.
@@
Notice that eq. 30 is precisely what our terms dependent on $\bm{ww}$ tell us, albeit projected onto the $\bm{ww}$-plane. This confirms that if we were to have a Gaussian profile, the total envelope $\bm{\Psi}+\bm{\Phi}$ must obey eq. 19. We can now turn our attention to the terms independent of $\bm{w}$. Note that if we add eq. 135 to our terms, we end up with:
$\frac{\mathrm{dln}(\alpha(\tau)A^{(0)}(\tau))}{\mathrm{d}\tau}+\frac{\mathrm{d}K_{\mu}}{\mathrm{d}\tau}\frac{\partial\hat{\bm{e}}}{\partial K_{\mu}}\cdot\hat{\bm{e}}^{*}+\hat{\bm{e}}^{*}\cdot\frac{\partial\bm{D}}{\partial K_{\mu}}\cdot\frac{\partial\hat{\bm{e}}}{\partial r_{\mu}}+\frac{1}{2}(\bm{\Psi}+\bm{\Phi}):\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H=0.$ (35)